# Impact of Large Language Model Assistance on Patients Reading Clinical Notes: A Mixed-Methods Study

### v1

@@ Abstract
Patients derive numerous benefits from reading their clinical notes, including an increased sense of control over their health and improved understanding of their care plan. However, complex medical concepts and jargon within clinical notes hinder patient comprehension and may lead to anxiety. We developed a patient-facing tool to make clinical notes more readable, leveraging large language models (LLMs) to simplify, extract information from, and add context to notes. We prompt engineered GPT-4 to perform these augmentation tasks on real clinical notes donated by breast cancer survivors and synthetic notes generated by a clinician, a total of 12 notes with 3868 words. In June 2023, 200 female-identifying US-based participants were randomly assigned three clinical notes with varying levels of augmentations using our tool. Participants answered questions about each note, evaluating their understanding of follow-up actions and self-reported confidence. We found that augmentations were associated with a significant increase in action understanding score (0.63 $\pm$ 0.04 for select augmentations, compared to 0.54 $\pm$ 0.02 for the control) with p=0.002. In-depth interviews of self-identifying breast cancer patients (N=7) were also conducted via video conferencing. Augmentations, especially definitions, elicited positive responses among the seven participants, with some concerns about relying on LLMs. Augmentations were evaluated for errors by clinicians, and we found misleading errors occur, with errors more common in real donated notes than synthetic notes, illustrating the importance of carefully written clinical notes. Augmentations improve some but not all readability metrics. This work demonstrates the potential of LLMs to improve patients’ experience with clinical notes at a lower burden to clinicians. However, having a human in the loop is important to correct potential model errors.
Introduction
Access to clinical documentation can be empowering for patients [1]. Patients have reported that reading their own notes helps them feel in control of their care and better remember the care plan [2]. Clinical notes can be challenging to understand, however. [3, 4]. One study found that only 24% of patients could comprehend their own discharge summaries [5]. A prior study of oncology notes showed that simplifying notes can improve patient understanding [6], and after providing patients access to their own notes, one hospital found that oncologists generated notes with increased readability [7]. There have also been additional efforts to improve readability of notes for patients [8, 9]. However, manual interventions are time-intensive and therefore difficult to scale [10, 11]. Artificial intelligence offers an opportunity to enhance clinical note readability. Given its potential impact, there has been significant existing work on biomedical text simplification [12]. Over the last decade, methods have largely consisted of (i) rule-based methods [12], which are brittle and difficult to generalize, and (ii) deep learning methods that have suffered from operating in a low-resource setting with a lack of health data. Recent works have begun to assess these solutions, primarily focusing on preliminary assessments with clinical raters [13, 14, 15, 16, 17].
Contributions:
The core questions we seek to answer are (1) can AI augmentations enhance patients’ experience with their clinical notes, and (2) are there negative impacts from these augmentations? We conduct a thorough analysis of potential impacts on patients and patient perspectives on large language model (LLM) use in medicine, with a focus on breast cancer, where comprehension can be particularly challenging [18, 19, 20]. We present an end-to-end LLM-enabled tool for improving patient clinical note comprehension. We also present generalizable methods for evaluating these tools using both quantitative and qualitative methods, and demonstrate results from error analysis, in-depth interviews from seven breast cancer patients, and survey responses from 200 participants.
Methods
Augmentations of Clinical Notes with GPT-4
Based on a literature review, we propose five types of augmentations to clinical notes to support patients: recognizing and defining medical terms in clinical notes (definitions), simplifying the original clinical note into plain English (simplification), generating common questions and answers based on the clinical note (FAQ), extracting key information from the clinical note (key information) and finally extracting any action items that the patient needs to do (to-do list). Each of the augmentations was obtained by passing a clinical note with a natural language instruction tailored for each augmentation as an input prompt to the LLM GPT-4 [21], as illustrated in Fig. 1. Our analysis drew from 12 clinical notes, containing 3868 words. Eight real notes were de-identified and donated by interview study participant, and were supplemented by four synthetic notes. The exact instructions and prompts used to obtain the augmentations can be found in the supplementary material (Fig. 6 and Section A). We employed prompt engineering strategies specific to each augmentation, primarily focusing on zero-shot prompting and prompt chaining [22]. These augmentations were then integrated into a user-friendly interface (Fig. 2, Supplementary Fig. 5).
Figure 1: Overview of mixed methods evaluation of LLM augmentations for clinical notes. This evaluation was conducted based on both synthetic and real notes, and two sets of augmentations were studied: select and all. Select included simplification and definition augmentations, and All included these as well as FAQ, Key Information, and To-Do List augmentations. Augmentation evaluation was conducted by mixed methods studies with participants, clinical review, and automated analyses. Several of these evaluations focused on identifying errors.
@@ Methods
Figure 2: Overview of the user interface, containing (from left to right) an option to toggle between the original clinical note and simplified augmentation, an optional pane with the FAQ, key information, and to-do list augmentations, and finally, an optional definitions augmentation pane.
Survey Study Design
@@
To test these augmentations’ effects on patient understanding, we first conducted a randomized web-based survey with generally healthy participants. We recruited 200 participants using the Prolific crowdsourcing platform with the following inclusion criteria: female-identifying, over 40 years of age, fluent in English. Participants first completed a survey about their demographics and a health literacy test, then viewed three notes. Study participants were randomized to either observe three synthetic or three real notes, which had been donated and de-identified. Notes were randomly selected from a pool of four synthetic and six real notes.
@@
Per note, augmentation, and participant, we obtained an action understanding score in the range $[0,1]$, a self-reported confidence score about understanding the care of the patient in $[0,3]$ (self-comprehension), and a self-reported confidence score about understanding the next steps of the patient in $[0,3]$ (self-confidence). At the conclusion of the study, we requested a list of augmentations they would like to use again, and participants additionally described what they used to answer the multiple-choice questions in a free-form text response. A clinician research partner reviewed all augmentations to check for false or misleading information. Participants were shown a disclaimer at the start of the study, warning that some content was generated using LLMs. Information identified as misleading or false was explicitly listed and corrected at the end of the study to avoid misleading participants. The study took approximately 20 minutes and participants were compensated with \$3.75. This study (protocol #E-5002) was determined exempt from IRB approval by the MIT Committee On the Use of Humans as Experimental Subjects.
Interview Study Design
@@
To evaluate the impact of our proposed clinical note augmentations and supplement the survey with more detailed feedback, we conducted in-depth interviews with patients with a history of breast cancer to explore patient perspectives. We recruited participants (N=7) who self-identified as breast cancer patients by advertising via the crowdsourcing platform Prolific and the Center for Clinical and Translational Research at MIT. Details about participants’ background is reported in Table 1. Interview studies were conducted virtually with two study team members present, including one clinical partner, and they followed a semi-structured interview protocol consisting of exploring our proposed augmentations interface and answering a series of questions about the augmentations (found in the supplementary material, Section E), including a Likert rating for each augmentation. This study (protocol #2208000725) was approved by the MIT Committee On the Use of Humans as Experimental Subjects.
Error Analysis of Augmentations
@@
Augmentations were manually reviewed for misleading or false information by three clinician research partners (BDL, TU, CER). Once validated, they were sorted into error categories (“Omission”, “Not Simplified”, “Incorrect Reasoning”, “Incorrect Fact”, “Misleading Phrasing”, “Awkward Phrasing”, “Hallucination”, “Timing”, “Brand Name”, “Unnecessary”, “Clinician Preference”, “Assumed Context”). Please see Supplemental Fig. H.
@@
In addition to manual review, we conducted an automated analysis of the augmentations. We report the Flesch-Kincaid Grade [23, 24] and the percentage of words in a dictionary (Carnegie Mellon Pronouncing Dictionary Corpus Reader [^1] https://www.nltk.org/_modules/nltk/corpus/reader/cmudict.html), to compare the acronyms and jargon in the original versus augmented notes.
Statistical Analysis
@@
To obtain 95% confidence intervals for mean estimates of outcomes, we use the method from Morey (2008) to correct for inter-subject variability [25]. To compare outcome means between two conditions, we use the pairwise Wilcoxon Signed Rank test to obtain p-values. We correct for multiple hypothesis testing with the Benjamin-Hochberg adjustment procedure with $\alpha=0.05$. All analyses were performed in R version 4.1.1 with the stats package.
Results
Study Population
@@ Results
For our survey study, we were able to recruit 200 participants through Prolific, many of whom had accessed the health portal, had adequate health literacy, and agreed that AI can have a beneficial impact in medicine. Most participants in both components of the study identified as female, and the majority of participants identified as white and non-Hispanic, Latino, or Spanish origin. The full demographic breakdown of the participants is in Table 1.

### v2

@@ Abstract
Large language models (LLMs) have immense potential to make information more accessible, particularly in medicine, where complex medical jargon can hinder patient comprehension of clinical notes. We developed a patient-facing tool using LLMs to make clinical notes more readable by simplifying, extracting information from, and adding context to the notes. We piloted the tool with clinical notes donated by patients with a history of breast cancer and synthetic notes from a clinician. Participants (N=200, healthy, female-identifying patients) were randomly assigned three clinical notes in our tool with varying levels of augmentations and answered quantitative and qualitative questions evaluating their understanding of follow-up actions. Augmentations significantly increased their quantitative understanding scores. In-depth interviews were conducted with participants (N=7, patients with a history of breast cancer), revealing both positive sentiments about the augmentations and concerns about AI. We also performed a qualitative clinician-driven analysis of the model’s error modes.
1 Introduction
@@
Figure 1: Our user interface to assist in the comprehension of clinical notes, containing (from left to right) an option to toggle between the original clinical note and simplified augmentation, an optional pane with the FAQ, key information, and to-do list augmentations, and finally, an optional definitions augmentation pane. Headings have been shown in the same formatting as visualizations in the rest of the paper.
The rapid advancement and improved performance of natural language processing could reduce barriers in communication and make information accessible to wider populations. Language models have the potential to help personalize and tailor the level and format of messaging to a variety of needs, e.g., to facilitate new avenues for public interest communication. Medicine, and public health more broadly, could particularly stand to benefit from new innovations in this space to facilitate improved patient-provider interaction [5].
The information needs of patients are not well addressed by existing clinical workflows. In appointments, clinicians frequently use unexplained medical terminology [32], which patients have trouble understanding [17, 41]. This information barrier has large implications; over a third of patients do not remember basic information about new medications [46]. In response, patients’ rights advocates have pushed for increased transparency by enabling patients to read and access their own clinical notes, which act as physician-authored summaries of an encounter. Reviewing clinical notes after an appointment can enable patients to better understand their medical conditions and improve care plan adherence [42]. In 2021, the $\textrm{21}^{st}$ Century Cares Act mandated that providers make clinical notes digitally available to all patients [9, 39].
Unfortunately, the clinical notes in electronic health records are written for the audiences of the clinicians themselves, other clinicians involved in care, and administrative, billing, and insurance stakeholders [11]. As patients are not the target audience, clinical notes are often filled with medical jargon that can be prohibitively difficult to understand. A recent study found it to be the most frequently reported barrier to reading clinical notes [19]. Another study found that improving the coherence of notes improved patient recall, but the intervention was too time consuming to be reasonably scalable [45].
In this work, we studied whether large language models (LLMs) can help bridge this gap and make clinical notes more accessible. We focused on breast cancer, where comprehension can be particularly challenging, and the corresponding benefits of improved health literacy are significant [42, 2, 39]. In particular, patients with cancer note a mismatch between the desired and actual levels of involvement in care. The core questions we sought to investigate were (1) whether and how AI augmentations enhance patients’ experience with their clinical notes, and (2) whether there are negative impacts from these augmentations. Our contributions are as follows:
  • An end-to-end tool for improving clinical note comprehension, with five different LLM-enabled augmentations (see Figure 1)
  • A quantitative survey study (N=200) on patient comprehension in which participants were exposed to three different levels of augmentations
  • In-depth interviews and qualitative analysis of seven patients with a history of breast cancer on their opinions and concerns about the interface and LLM augmentations
  • An analysis of the types of errors induced by the LLM augmentations, as rated by three clinical partners
2 Related Work
Given the potential impact of increased accessibility to clinical notes, there have been several studies investigating the benefits of the status quo, current concerns, and possible mitigations. Several works have explored the existing impact of accessible clinical notes through interviews with patients and their care partners; benefits have included improved memory of the care plan, a better understanding of the medical condition and potential side effects, reassurance, and the ability for care teams to support patients remotely [43, 20, 14, 18, 12]. Further, the populations likelier to report benefits from reviewing their clinical notes are patients who are less educated, nonwhite, older, or with English as their nonprimary language [47]. However, many patients have made recommendations on how to improve the writing of clinical notes, citing difficulties in note comprehension due to the jargon in notes and not knowing where to focus attention while reading [19, 27, 25].
Given these known flaws in the current state of clinical notes, there have been suggestions for clinicians on how to improve note readability [24, 26]. After access to notes, one study found that oncologists were writing slightly longer notes with increased readability, as measured by automatic scores [37]. Others found that manual interventions, such as adding definitions, improving coherence of text, and expanding acronyms, did improve patient understanding, as measured by patient questionnaires [45, 29, 22]. Unfortunately, the interventions were manual and difficult to scale, and the burden of documentation is already considered a leading cause of physician burnout [38].
In response, the field of clinical natural language processing has pursued how to automate the simplification of clinical and biomedical text [34]. Until a few years ago, methods have largely consisted of (i) rule-based methods, which are brittle and difficult to generalize, and (ii) deep learning methods that have suffered from operating in a low-resource setting with a lack of health data. These had seen limited success in clinical text [34]. More recently, researchers have started exploring large language models as another avenue in preliminary studies for text simplification, in medicine and in the general domain [21]; other avenues have also been explored for patient communication including chatbots and question answering [8, 3]. In the general domain, text simplification has often focused on shorter documents, e.g. sentence-level [21, 49]. Clinically, LLMs have been found to improve the readability of radiology reports as measured by automated metrics, though these have been found to only loosely correlate with patient comprehension [28, 50]. Several works have found positive benefits from LLM text simplification in smaller user studies, as measured by patient perception [31, 40]. In contrast, we conducted a much larger, patient-centric study, that complements qualitative perception with quantitative measures of patient comprehension of realistic medical action items for the patient.
Much of the benchmarking of LLMs in medicine has focused on multiple choice question answering. Meanwhile, while difficult to qualify, large language models still have flaws when conducting open-ended generation in medicine. Several papers have found the introduction of unsupported facts, sometimes referred to as hallucinations, as a nontrivial problem in medical summarization [7]. In one study of open-ended medical question answering, physicians rated responses across axes such as bias and medical knowledge [44]. In our study, we studied a much more constrained scenario than open-ended question answering, as there is long context (the clinical note) that the response should be conditioned on, and we described a taxonomy of LLM errors for this longer-form clinical setting, based on thematic analysis.
3 Clinical Notes Augmentations
3.1 Augmentation Types
We propose five types of augmentations to clinical notes to support patients. Each of the augmentations was obtained by passing a clinical note with a natural language instruction tailored for each augmentation as an input prompt to the LLM GPT-4 (specifically version GPT-4-0314) [35]. We employed prompt engineering strategies specific to each augmentation, primarily focusing on zero-shot prompting and prompt chaining [48].
  • Definitions: recognizing and defining medical terms in clinical notes. The prompt was to “Make a list of the medical words that appear in this clinical note. Write each word as it appears in the note without modification. Define each of these terms. Definitions should be 1-2 sentences in simple English.”
  • Simplification: translating the original clinical note into layperson language. The prompt was to “Rewrite this clinical note in simple English. Do not remove names of medications or diagnoses.”
  • FAQ: generating common questions and answers based on the clinical note. The prompt was to “Make a list of 7 common questions a patient might ask their doctor after reading this clinical note. Answer the questions in 2-3 sentences using simple English.”
  • Key Information: extracting key information from the clinical note. The prompt was to “Make a list of the medications mentioned in this clinical note. Make a list of the upcoming appointments explicitly mentioned in this clinical note. If the note does not mention appointments, say ‘No Appointments’”
  • To-do List: extracting any action items that the patient needs to remember. The prompt was to “Make a todo list of tasks the patient is responsible for before their next appointment based on the clinical note. Tasks need to be specific and actionable. Tasks should not be recurring events. If there are no such tasks mentioned in the note, say ’No Tasks’”
Once we generated each of these augmentations, we integrated them into a user interface for patients to interact with to improve their experience reading clinical notes.
3.2 User Interface
We developed a user interface, implemented as a web app, to integrate the augmentations presented in Section 3.1 (Figure 1). The interface consists of three panels, each containing different augmentations. In the first panel, the reader has access to a tab containing the original clinical note and a second tab containing the generated simplified note. On the original note, each medical term is highlighted in bold green and underlined to indicate that the user can click on it. When a user clicks on a medical term, the page will automatically scroll to the corresponding definition which is available in the third panel. The second panel consists of three tabs for the augmentations FAQ, Key Information, and the To-Do List. For FAQ, the reader can expand on each question to see its answer. In Figure 3, we show explicit examples of the augmentations generated for one of the clinical notes used in our work.
Figure 2: Overview of mixed methods evaluation of LLM augmentations for clinical notes in our work. This evaluation was conducted based on both synthetic and real notes (left of figure), and two sets of augmentations were studied: select and all. Select included simplification and definition augmentations, and All included these as well as FAQ, Key Information, and To-Do List augmentations. Augmentation evaluation was conducted using mixed methods (see list under Augmentation Analysis). Researchers on the team analyzed the results of the first several methods, and clinical research partners focused on clinical review (person with + icon). Readability was analyzed automatically. We used methods including both surveys (pen icon) and interviews (chat bubble icon), as well as clinical review (person with + icon) and automated analyses (computer monitor icon). Several of these evaluations focused on identifying errors (those highlighted in last column).
Figure 3: Examples of the augmentations considered in our interface using a sample excerpt from a synthetic clinical note. This includes the augmentations of simplification, definitions, frequently asked questions, key information, and to-do list.
4 Experimental Setup
4.1 Clinical Notes Collection
To evaluate our interface, we needed a set of clinical notes for experimentation. We relied on two sources of clinical notes: synthetic notes written by a clinical research partner and real notes donated by patients with a history of breast cancer. The clinical research partner created four synthetic notes that mirror real clinical notes. These notes were inspired by the clinician research partner’s professional experience, but do not contain information about any of their patients. The difficulty of a clinical note is related to several factors, such as the extent of abbreviation usage and whether note-taking style sentences are used instead of full grammatical sentences. These synthetic notes were written at an intermediate difficulty, in the style of a clinician who is not writing for the purpose of patient comprehension.
We obtained eight real notes that were de-identified and donated by interview study participants (recruitment details can be found in Section 4.3). The participants were self-identified patients with a history of breast cancer. Participants were told to look for clinical notes containing free text describing an appointment during their breast cancer treatment. Suitable notes might contain sections labeled “Impression”, “Assessment & Plan” or “Plan”. We received written consent from participants before they donated their clinical notes and additional written consent before participating in the video interview. All participants were given a verbal overview of the purpose and risks listed in the consent forms, and were given an opportunity to ask questions to the research team before signing. This data collection process was approved by our Institutional Review Board (IRB).
In total, our analysis drew from 12 clinical notes, each containing an average of 322 words. We focused on this limited set of notes because the clinician research partner determined the corpus to represent a wide variety of clinical scenarios and note-writing styles, and because of the intensive time required to curate synthetic notes and collect and de-identify donated notes.
4.2 Survey Study Design
@@
To test these augmentations’ effects on patient understanding, we first conducted a randomized web-based survey with healthy participants. We recruited participants using the Prolific crowdsourcing platform with the following inclusion criteria: female-identifying, over 40 years of age, fluent in English. Participants first completed a survey about their demographics and a health literacy test (Confident with Forms [10]), then viewed three notes. Study participants were randomized to either observe synthetic or real (donated, de-identified) notes. Notes were randomly selected from a pool of four synthetic and six real notes.
@@
Per note, augmentation, and participant, we obtained an action understanding score in the range $[0,1]$, a self-reported confidence score about understanding the care of the patient in $[0,3]$ (self-comprehension), and a self-reported confidence score about understanding the next steps of the patient in $[0,3]$ (self-confidence). At the conclusion of the study, we requested a list of augmentations they would like to use again, and participants additionally described what they used to answer the multiple-choice questions in a free-form text response. A clinician research partner reviewed all augmentations to check for false or misleading information. Participants were shown a disclaimer at the start of the study, warning that some content was generated using LLMs. Information identified as misleading or false was explicitly listed and corrected at the end of the study to avoid misleading participants. The study took approximately 20 minutes and participants were compensated with \$3.75. Full details of the survey study can be found in Appendix E. This study was determined exempt from IRB approval by our institutional IRB.
4.3 Interview Study Design
@@
To evaluate the impact of our proposed clinical note augmentations and supplement the survey with more detailed feedback, we also conducted in-depth interviews to explore patient perspectives. We recruited participants who self-identified as patients with a history of breast cancer by advertising via the crowdsourcing platform Prolific and a clinical center at our institution. Details about participants’ backgrounds is reported in Table 1. Interview studies were conducted virtually with two study team members present, including one clinical partner, following a semi-structured interview protocol consisting of exploring our proposed augmentations interface and answering a series of questions about the augmentations, including a Likert rating for each augmentation.
One researcher conducted all of the interviews virtually. A clinical research partner was present for all interviews to answer general safety-related medical questions as needed. All sessions lasted one hour and were audio- and video-recorded via a teleconference platform, then transcribed manually. Both researchers introduced themselves at the start of the interview and reviewed the study goals. Participants were asked to answer demographic questions. A semi-structured interview protocol was developed based on the initial survey findings. Participants were asked about their prior experience reading clinical notes, and then were given a customized link to the user interface where one of their donated clinical notes and its augmentations had been preloaded. They were asked to think aloud as they used the tool, and specifically were asked what they liked and disliked, whether anything about the augmentation worried them and if they identified any errors, and finally to rate on a Likert scale from 1-5 how much the augmentation assisted their reading experience. Full details of the interview protocol and questions can be found in Appendix B. This study was approved by our IRB.
Two researchers used DeDoose software and Braun and Clarke’s methods for thematic analysis for qualitative analysis of the transcribed interviews [6]. One researcher first analyzed the interviews to derive major themes for an initial codebook, which was reviewed by the second researcher. Both researchers independently coded all of the interviews and iteratively updated the codebook and reconciled discrepant codes until consensus was reached.
4.4 Error Analysis of Augmentations
@@
All augmentations underwent manual review by three clinician research partners to identify misleading or false information. One clinician research partner conducted an initial examination of all notes and augmentations, documenting any errors found along with their rationale for categorizing them as errors. The other two clinicians then validated these findings, to ensure consistency in our findings. Once evaluated by all three clinician research partners, they were sorted into error categories (“Omission”, “Not Simplified”, “Incorrect Reasoning”, “Incorrect Fact”, “Misleading Phrasing”, “Awkward Phrasing”, “Hallucination”, “Timing”, “Brand Name”, “Unnecessary”, “Clinician Preference”, “Assumed Context”) by multiple coauthors, including the clinician research partners and non-clinical research partners. We worked together to reach consensus assigning categories to each error found. The full categories are described in Appendix D.
@@
In addition to manual review, we also conducted an automated analysis of the augmentations. We reported the Flesch-Kincaid Grade [36, 23] and the percentage of words in a dictionary (Carnegie Mellon Pronouncing Dictionary Corpus Reader [^1] https://www.nltk.org/_modules/nltk/corpus/reader/cmudict.html), to compare the acronyms and jargon in the original versus augmented notes.
4.5 Statistical Analysis
@@
To obtain 95% confidence intervals for mean estimates of outcomes, we used a method to correct for inter-subject variability [33]. To compare outcome means between two conditions, we used the pairwise Wilcoxon Signed Rank test to obtain p-values. We corrected for multiple hypothesis testing with the Benjamin-Hochberg adjustment procedure with $\alpha=0.05$. We also conducted subgroup analyses, such as for those who never accessed a health portal, cancer patients and health professionals, those with an associate degree or less, and those who were not White race, to see whether overall trends held. All analyses were performed in R version 4.1.1 with the stats package.
5 Results
5.1 Study Population
@@
For our survey study, we recruited 200 participants through Prolific, many of whom had accessed a health portal, had adequate health literacy, and agreed that AI can have a beneficial impact in medicine. Most participants in both components of the study identified as female, and the majority of participants identified as white and non-Hispanic, Latino, or Spanish origin. The full demographic breakdown of the participants is in Table 1.